OpenStudy (emilyjones284):

please help..ill give a medal (:

OpenStudy (emilyjones284):

\[2x=\sqrt{20+2x}\]

OpenStudy (anonymous):

Start by squaring both sides. This eliminates the radical: \[4x^2 = 20 + 2x\] Subtract the 20 and the 2x from both sides to get a quadratic equation: \[4x^2 -2x -20=0\] From here you could try factoring or plugging in to the quadratic formula. If you need more help with solving this let me know.

jhonyy9 (jhonyy9):

like first step you can factorizing out 2 and divide both sides by 2 and so will get 2x^2 -x -10 = 0

OpenStudy (emilyjones284):

ohhhhh ok thank you!